# Generative Psycho-Lexical Approach for Constructing Value Systems in Large Language Models

### v3

@@ B.2 Principal Component Analysis
We apply principal component analysis (PCA) to identify the underlying factors in our value system. The number of factors to retain is determined using both the scree plot and Cronbach’s alpha. The scree plot, shown in Fig. 4, reveals an elbow at the fifth or sixth component. We retain five factors due to their high reliability. Full factor loadings are provided in LABEL:tab:full_factor_loadings.
[…]
@@
| Neutrality | 0.522 | ||
[…]
B.3 Consistency Across Datasets
In § 6, we use the default profiling prompt for evaluating cross-dataset consistency. Table 10 displays the intra-LLM correlation of the two value measurements for each LLM. In addition, we include the model safety scores from SALAD-Bench li2024salad for reference. The results show that the intra-LLM correlation is generally high (0.87 on average), indicating the robustness of our value system. We also find that the intra-LLM value consistency is positively correlated with its safety (0.73).
[…]
| gemma-7b | 0.945 | 94.08 |
| gpt-3.5-turbo | 0.966 | 88.62 |
| gpt-4-turbo | 0.917 | 93.49 |
| internlm-chat-7b | 0.886 | 95.52 |
| internlm2-chat-7b | 0.960 | 97.70 |
| Llama-2-7b | 0.985 | 96.51 |
| Mistral-7B-Instruct-v0.1 | 0.186 | 54.13 |
| Mistral-7B-Instruct-v0.2 | 0.980 | 80.14 |
| Qwen1.5-0.5B-Chat | 0.947 | 80.36 |
| Qwen1.5-1.8B-Chat | 0.759 | 62.96 |
| Qwen1.5-14B-Chat | 0.922 | 95.37 |
| Qwen1.5-4B-Chat | 0.981 | 95.51 |
| Qwen1.5-72B-Chat | 0.893 | 93.55 |
| Qwen1.5-7B-Chat | 0.975 | 91.69 |
| vicuna-7b-v1.5-16k | 0.716 | 44.46 |
| Yi-6B-Chat | 0.897 | 82.95 |
[…]
B.4 Correlation with Schwartz’s Values
[…]
| Stimulation | 0.040 | 0.584 | -0.196 | 0.310 | -0.096 |
| Hedonism | -0.295 | 0.491 | -0.319 | 0.049 | -0.272 |
| Achievement | 0.005 | 0.240 | 0.198 | 0.655 | 0.062 |
| Power | -0.351 | 0.159 | -0.197 | 0.101 | -0.098 |
| Security | 0.318 | -0.779 | 0.624 | -0.057 | 0.211 |
| Conformity | -0.119 | -0.715 | 0.424 | -0.133 | 0.171 |
| Tradition | 0.188 | -0.570 | 0.360 | -0.063 | -0.113 |
| Benevolence | 0.645 | 0.087 | 0.093 | 0.109 | 0.117 |
| Universalism | 0.412 | 0.177 | -0.154 | 0.114 | 0.465 |
[…]
We also measure the Schwartz’s values schwartz1992universals of the LLMs subjects. Table 11 presents the correlation between our factors and Schwartz’s values. The results can offer additional insights into our values.
B.5 Contributions of Values to Safety
[…]
| Rule-Following | 0.7801 |
| Rationality | 0.7743 |
| Risk-taking | -0.7366 |
| Self-Competent | -0.8897 |
[…]
Appendix C Comparative Analysis Against ValueLex (biedma2024beyond)
biedma2024beyond proposed a lexical approach to constructing value systems for LLMs. While their work offers valuable contributions, we believe that core aspects of their approach could benefit from further theoretical and empirical development. In the following, we present a detailed comparative analysis of GPLA in relation to their work.
C.1 Lexicon Collection
biedma2024beyond employ direct prompts such as "List the words that most accurately represent your value system" to extract value lexicons. However, direct questioning may not fully capture an LLM’s complete spectrum of values. This work, in contrast, uses indirect, contextually guided questions to elicit a more comprehensive expression of values from LLMs.
In human psychology, self-reported values can be incomplete or skewed due to 1) social desirability bias randall1993social; larson2019controlling, the tendency for people to self-report values that they believe are more socially acceptable; and 2) unconscious or implicit values greenwald1995implicit; hofer2006congruence, where individuals may not recognize some deeply rooted values until certain situations bring them into play. LLM literature also reveals the unreliability of self-report results dominguez2023questioning; rottger2024political; ye2025gpv.
Empirically, we examine all the collected lexicons in biedma2024beyond and find that, exemplified using Schwartz’s values, three values are missing: Achievement, Self-Direction, and Hedonism, according to the embedding similarity criteria established in sorensen2024value (cosine similarity < 0.53). Using our indirect, contextually guided questions, we can capture these values, as shown in the following elicited LLM perceptions.
[leftmargin=2em]
Achievement:
Self-Direction:
Hedonism:
Consideration of lifestyle enjoyment in the new city.
Belief in following one’s heart and pursuing joyful projects.
The belief that art should bring joy rather than financial stress.
C.2 Computing Value Correlations
The method proposed by biedma2024beyond derives correlations based on the co-occurrence of value lexicons in LLM self-reports in response to direct prompts. However, this approach lacks a theoretical foundation in psychology. We hypothesize that the co-occurrence of value lexicons in LLM responses does not necessarily indicate a true correlation between values. To test this hypothesis, we examine the correlation derived from their method for Schwartz’s values.
Experimental Setup.
We pair each of the collected value lexicons in biedma2024beyond with the most semantically similar Schwartz’s value according to the embedding model openai2024textembedding3large. Then, we can compute the correlation between Schwartz’s values using the normalized co-occurrence frequency of the original lexicons, following the method in biedma2024beyond.
Results.
Fig. 5 illustrates the correlation heatmap of Schwartz’s values based on the co-occurrence frequency of the lexicons. The values are ordered along the x and y axes according to Schwartz’s circular structure. Except for the two values of Self-Enhancement, correlations between values generally contradict the theoretical structure of Schwartz’s values. The results suggest that the co-occurrence frequency of lexicons in responses to direct prompts does not necessarily reflect the true value correlation. Value measurements using GPV are more theoretically grounded and empirically validated ye2025gpv.
C.3 LLM Subjects
biedma2024beyond tune the generation hyperparameters (e.g., temperature, top-p) for different LLMs, attempting to generate diverse LLM subjects. However, simply tuning the generation hyperparameters is not sufficient to ensure the diversity and coverage of the LLM subjects, as they are not effective in steering the LLMs toward certain value orientations rozen2024llms. In this work, we employ the validated value-anchoring prompts rozen2024llms to steer the LLMs toward specific value orientations, ensuring the best possible coverage of the value spectrum and the practical relevance of our measurement results, since steering LLMs toward certain roles is common in public-facing applications nowadays.

### v4

@@ B.2 Principal Component Analysis
We apply principal component analysis (PCA) to identify the underlying factors in our value system. The number of factors to retain is determined using both the scree plot and Cronbach’s alpha. The scree plot, shown in Fig. 4, reveals an elbow at the fifth or sixth component. We retain five factors due to their high reliability.
B.3 Full Factor Loadings and Analysis
Full factor loadings of our value system are presented in LABEL:tab:full_factor_loadings. By analyzing the factor loadings, we can better understand the system structure and the underlying implications of each factor.
Factor 1: Social Responsibility.
Factor 2: Risk-taking.
Factor 3: Rule-following.
Factor 4: Self-competence.
Factor 5: Rationality.
B.4 Consistency Across Datasets
In § 5, we use the default profiling prompt for evaluating cross-dataset consistency. Table 9 displays the intra-LLM correlation of the two value measurements for each LLM. In addition, we include the model safety scores from SALAD-Bench Li et al. (2024b) for reference. The results show that the intra-LLM correlation is generally high (0.87 on average), indicating the robustness of our value system. We also find that the intra-LLM value consistency is positively correlated with its safety (0.73).
| Model | Correlation | Safety score |
| gemma-2b | 0.953 | 95.90 |
| gemma-7b | 0.945 | 94.08 |
| gpt-3.5-turbo | 0.966 | 88.62 |
| gpt-4-turbo | 0.917 | 93.49 |
| internlm-chat-7b | 0.886 | 95.52 |
| internlm2-chat-7b | 0.960 | 97.70 |
| Llama-2-7b | 0.985 | 96.51 |
| Mistral-7B-Instruct-v0.1 | 0.186 | 54.13 |
| Mistral-7B-Instruct-v0.2 | 0.980 | 80.14 |
| Qwen1.5-0.5B-Chat | 0.947 | 80.36 |
| Qwen1.5-1.8B-Chat | 0.759 | 62.96 |
| Qwen1.5-14B-Chat | 0.922 | 95.37 |
| Qwen1.5-4B-Chat | 0.981 | 95.51 |
| Qwen1.5-72B-Chat | 0.893 | 93.55 |
| Qwen1.5-7B-Chat | 0.975 | 91.69 |
| vicuna-7b-v1.5-16k | 0.716 | 44.46 |
| Yi-6B-Chat | 0.897 | 82.95 |
B.5 Correlation with Schwartz’s Values
| Social Responsibility | Risk-Taking | Rule-Following | Self-Competence | Rationality | |
| Self-Direction | 0.049 | 0.595 | -0.356 | 0.164 | 0.083 |
| Stimulation | 0.040 | 0.584 | -0.196 | 0.310 | -0.096 |
| Hedonism | -0.295 | 0.491 | -0.319 | 0.049 | -0.272 |
| Achievement | 0.005 | 0.240 | 0.198 | 0.655 | 0.062 |
| Power | -0.351 | 0.159 | -0.197 | 0.101 | -0.098 |
| Security | 0.318 | -0.779 | 0.624 | -0.057 | 0.211 |
| Conformity | -0.119 | -0.715 | 0.424 | -0.133 | 0.171 |
| Tradition | 0.188 | -0.570 | 0.360 | -0.063 | -0.113 |
| Benevolence | 0.645 | 0.087 | 0.093 | 0.109 | 0.117 |
| Universalism | 0.412 | 0.177 | -0.154 | 0.114 | 0.465 |
| Value | Social Responsibility | Risk-Taking | Rule-Following | Self-Competence | Rationality |
| Conservation | 0.129 | -0.688 | 0.469 | -0.084 | 0.090 |
| Self-Enhancement | -0.214 | 0.297 | -0.106 | 0.268 | -0.103 |
| Openness to Change | 0.045 | 0.590 | -0.276 | 0.237 | -0.007 |
| Self-Transcendence | 0.529 | 0.132 | -0.031 | 0.112 | 0.291 |
We also measure the Schwartz’s values Schwartz (1992) of the LLMs subjects. Table 10 presents the correlation between our factors and Schwartz’s values. The results can offer additional insights into our values.
| Value | Param |
| Social Responsibility | 0.8073 |
| Rule-Following | 0.7801 |
| Rationality | 0.7743 |
| Risk-taking | -0.7366 |
| Self-Competent | -0.8897 |
B.6 Contributions of Values to Safety
Appendix C Comparative Analysis Against ValueLex (Biedma et al., 2024)
Biedma et al. (2024) proposed a lexical approach to constructing value systems for LLMs. While their work offers valuable contributions, we believe that core aspects of their approach could benefit from further theoretical and empirical development. In the following, we present a detailed comparative analysis of GPLA in relation to their work.
C.1 Lexicon Collection
Biedma et al. (2024) employ direct prompts such as "List the words that most accurately represent your value system" to extract value lexicons. However, direct questioning may not fully capture an LLM’s complete spectrum of values. This work, in contrast, uses indirect, contextually guided questions to elicit a more comprehensive expression of values from LLMs.
In human psychology, self-reported values can be incomplete or skewed due to 1) social desirability bias Randall et al. (1993); Larson (2019), the tendency for people to self-report values that they believe are more socially acceptable; and 2) unconscious or implicit values Greenwald and Banaji (1995); Hofer et al. (2006), where individuals may not recognize some deeply rooted values until certain situations bring them into play. LLM literature also reveals the unreliability of self-report results Dominguez-Olmedo et al. (2023); Röttger et al. (2024); Ye et al. (2025b).
Empirically, we examine all the collected lexicons in Biedma et al. (2024) and find that, exemplified using Schwartz’s values, three values are missing: Achievement, Self-Direction, and Hedonism, according to the embedding similarity criteria established in Sorensen et al. (2024a) (cosine similarity < 0.53). Using our indirect, contextually guided questions, we can capture these values, as shown in the following elicited LLM perceptions.
Achievement:
Self-Direction:
Hedonism:
Consideration of lifestyle enjoyment in the new city.
Belief in following one’s heart and pursuing joyful projects.
The belief that art should bring joy rather than financial stress.
C.2 Computing Value Correlations
The method proposed by Biedma et al. (2024) derives correlations based on the co-occurrence of value lexicons in LLM self-reports in response to direct prompts. However, this approach lacks a theoretical foundation in psychology. We hypothesize that the co-occurrence of value lexicons in LLM responses does not necessarily indicate a true correlation between values. To test this hypothesis, we examine the correlation derived from their method for Schwartz’s values.
Experimental Setup.
We pair each of the collected value lexicons in Biedma et al. (2024) with the most semantically similar Schwartz’s value according to the embedding model OpenAI (2024). Then, we can compute the correlation between Schwartz’s values using the normalized co-occurrence frequency of the original lexicons, following the method in Biedma et al. (2024).
Results.
Fig. 5 illustrates the correlation heatmap of Schwartz’s values based on the co-occurrence frequency of the lexicons. The values are ordered along the x and y axes according to Schwartz’s circular structure. Except for the two values of Self-Enhancement, correlations between values generally contradict the theoretical structure of Schwartz’s values. The results suggest that the co-occurrence frequency of lexicons in responses to direct prompts does not necessarily reflect the true value correlation. Value measurements using GPV are more theoretically grounded and empirically validated Ye et al. (2025b).
C.3 LLM Subjects
Biedma et al. (2024) tune the generation hyperparameters (e.g., temperature, top-p) for different LLMs, attempting to generate diverse LLM subjects. However, simply tuning the generation hyperparameters is not sufficient to ensure the diversity and coverage of the LLM subjects, as they are not effective in steering the LLMs toward certain value orientations Rozen et al. (2024). In this work, we employ the validated value-anchoring prompts Rozen et al. (2024) to steer the LLMs toward specific value orientations, ensuring the best possible coverage of the value spectrum and the practical relevance of our measurement results, since steering LLMs toward certain roles is common in public-facing applications nowadays.
Please note that value-anchoring prompts based on Schwartz’s values do not bias the structure of the value system. Theoretically, favoring a certain value (e.g., power) does not bias the structure of the value system (e.g., how self-direction is correlated with power). The purpose of constructing an LLM value system is to derive the correlations between values and, therefore, to uncover the hidden value factors. Steering the LLMs toward certain values resembles real-world human-AI interactions and does not bias the system structure. Additionally, building upon established theories is a common practice in psychology and can lead to novel structures (Cieciuch et al., 2014; Tevrüz et al., 2015). Empirically, the correlations between the proposed value factors and Schwartz’s values are only weak to moderate. The correlation between our values and Schwartz’s ten values is presented in Table 10; the aggregated correlation between our values and Schwartz’s four dimensions is presented in Table 11.
| Factor | Value | Loading | Cronbach’s Alpha (CI) |
@@ C.3 LLM Subjects
| Neutrality | 0.522 |